# Policy Search, Retrieval, and Composition via Task Similarity in Collaborative Agentic Systems

### v2

@@ Appendix A Extended Related Work
Appendix B Extending MOSAIC to Lifelong Reinforcement Learning
@@
where each $\beta$ is a scalar weight normalized via softmax. In this case, the initialization problem includes setting the weighting for previous task masks, the current task mask, and the newly acquired masks from other agents. One simple approach could simply be to extend Equation (8) to
@@
Appendix C Additional Ablation Studies
@@
We conducted a grid search on the number of samples per embedding $N$ (Equation (2)) and the size of the reference distribution $M$ (Equation (3)) to determine their effects on similarity computations and performance on the image sequence learning benchmark. Results shown in Figure 8 indicate no statistically significant changes from using different parameters of $N$ and $M$, on the resulting performance.
@@
Appendix D Additional Experiment Analysis
Figure 11 illustrates an example of the sequence-learning problem based on (Soltoggio et al. 2019). The agent must navigate a series of nodes labeled as H (home), W (wait state), D (decision state), and R (reward state). For the experiments in this paper, the wait state probability is set to zero, resulting in reduced complexity.
@@ Appendix D Additional Experiment Analysis
D.1 Further Analysis of Similarity and Converged Beta Parameters
@@
Appendix E Implementation Details
@@
Network Architecture Details
Critic: Linear (200, 1)
Actor: Linear (200, action)
Critic: Linear (256, 1)
Actor: Linear (256, action)

### v3

@@ Appendix B Extended Related Work
Appendix C Extending MOSAIC to Lifelong Reinforcement Learning
@@
where each $\beta$ is a scalar weight normalized via softmax. In this case, the initialization problem includes setting the weighting for previous task masks, the current task mask, and the newly acquired masks from other agents. One simple approach could simply be to extend Eq. (8) to
@@
Appendix D Additional Ablation Studies
@@
We conducted a grid search on the number of samples per embedding $N$ (Eq. (2)) and the size of the reference distribution $M$ (Eq. (3)) to determine their effects on similarity computations and performance on the image sequence learning benchmark. Results shown in Figure 8 indicate no statistically significant changes from using different parameters of $N$ and $M$, on the resulting performance.
@@
Appendix E Additional Experiment Analysis
Figure 11 illustrates an example of the sequence-learning problem based on (Soltoggio et al. 2023). The agent must navigate a series of nodes labeled as H (home), W (wait state), D (decision state), and R (reward state). For the experiments in this paper, the wait state probability is set to zero, resulting in reduced complexity.
@@ Appendix E Additional Experiment Analysis
E.1 Further Analysis of Similarity and Converged Beta Parameters
@@
Appendix F Implementation Details
@@
Network Architecture Details
Critic: Linear (200, 1)
Actor: Linear (200, action)
Critic: Linear (256, 1)
Actor: Linear (256, action)